OpenStudy (anonymous):

Find the slope of the tangent line to the ellipse given by x2 + 4y2 = 4 at the point (√2, -1/√2)

OpenStudy (anonymous):

implicit differentiation

OpenStudy (anonymous):

\[2x+8y \frac{dy}{dx}=0\]

OpenStudy (anonymous):

\[8y \frac{dy}{dx}=-2x\]

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{-2x}{8y}\]

OpenStudy (anonymous):

reduce and substitute x and y into derivative to get slope

OpenStudy (anonymous):

but when using the long way, it came out to be -1/2 and the answer should be 1/2 when using implicit differentiation

OpenStudy (anonymous):

save this it will help you a lot http://www.wolframalpha.com/

OpenStudy (anonymous):

I got -1/2 through implicit... make sure you subtract 2x from both sides to get the negative.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

see

OpenStudy (anonymous):

yeah, I got 1/2

OpenStudy (anonymous):

negative from derivative and negative in the y term make positive

OpenStudy (anonymous):

did that help you any

OpenStudy (anonymous):

???

OpenStudy (anonymous):

that is implicit differentiation, but when i do it in the other way, it came out to be an opposite sign answer